The length of the rug is 4 feet more than Its width. Find the dimensions of the rug if it's area is
x^2-8x+12

Answer by Boreal(15235)

length=x+4
width=x
area=x(x+4)=x^2+4x
But x^2+4x=x^2-8x+12
solving
12x=12, since x^2 cancel on each side
x=1 foot
x+4=5 feet
area is 5 square feet.
x^2-8x+12=1-8+12=5 sq feet.
The rug is 1 x 5 feet